M is the midpoint of XY. X has coordinates (2, 7)and M has coordinates (6, 1). Find the coordinates of Y.

Answer by midwood_trail(310)

M is the midpoint of XY. X has coordinates (2, 7)and M has coordinates (6, 1). Find the coordinates of Y.
Let (x,y) be the coordinates of Y. Then, by the midpoint formula:
midpoint = [(x1 + x2)/2 , (y1 + y2)/2]
We are given the midpoint to be the point (6, 1).
Let midpoint = (6, 1)
(6, 1) = [(x + 2)/2 , (y + 7)/2]
Do you see the point (6,1)?
To find the coordinates of the y point, we equate (x + 2)/2 to 6
and then we equate (y + 7)/2 to 1 and solve for x and y individually.
(x + 2)/2 = 6
x + 2 = 6 * 2
x + 2 = 12
x = 12 - 2
x = 10
=============
(y + 7)/2 = 1
y + 7 = 1 * 2
y + 7 = 2
y = 2 - 7
y = -5
The coordinates of point y are x = 10 and y = -5 written (10, -5).
Did you follow?